a) Does the sequence (3+(-1)n) converge to 2? Justify.

 

 b) Show that "\lim _{x\to \infty }\left(\frac{x-3}{x+1}\right)^x=e^{-4}"


c) Check whether the sequence fn (x) = "\frac{3x}{1+nx^2}" where x ∈ [2,∞ [ is uniformly 

convergent in [2,∞ [


1
Expert's answer
2022-05-25T15:21:32-0400

1) The sequence 

"\left(3+\left(-1\right)^n\right)"

doesn't converge to 2, because this sequence is oscillating between 2 and 4. Hence, it is divergent.


2) "\lim_{x\rightarrow\infty}{\left(\frac{x-3}{x+1}\right)^x}=\lim_{x\rightarrow\infty}{\left(\frac{x+1-4}{x+1}\right)^x}=" "\lim\limits_{x\rightarrow\infty}{\left(\left(1+\frac{-4}{x+1}\right)^\frac{x+1}{-4}\right)^{-\frac{4x}{x+1}}}="

"=\lim_{x\rightarrow\infty}{e^{-\frac{4}{1+\frac{1}{x}}}}= e^{-4}".


3) if x=2 then "f_n\left(2\right)=\frac{6}{1+4n}" and "\lim_{n\rightarrow\infty}{f_n\left(2\right)}=\lim_{n\rightarrow\infty}{\frac{6}{1+4n}}=0".


If x>2 then "\left|f_n\left(x\right)\right|=\left|\frac{3x}{1+nx^2}\right|\le\left|\frac{3x}{nx^2}\right|=\left|\frac{3}{nx}\right|" and "\lim_{n\rightarrow\infty}{\frac{3}{nx}}=0"

Conclution:

"f_n\left(x\right)\rightrightarrows0,\ x\in\left[2,+\infty\right["